OpenStudy (anonymous):

LAST QUESTION: Solve the system of equations y=x-6 x+y=-2

mathslover (mathslover):

\[\large{y=x-6}\] \[\large{x+y=-2}\] take first equation first

mathslover (mathslover):

that is we have y = x-6

mathslover (mathslover):

put this value of y in the second equation... I mean to say that : \[\large{x+y=-2}\] \[\large{x+(x-6)=-2}\]

mathslover (mathslover):

can you do that further @Tutu ? what do you get for solving for x

OpenStudy (anonymous):

yeahh umm i dont understand what you tryna say.

mathslover (mathslover):

oh k no problem

mathslover (mathslover):

I will explain u more

mathslover (mathslover):

given y = x-6 right?

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yeah sure right

mathslover (mathslover):

so the value of y is x-6

mathslover (mathslover):

right?

OpenStudy (anonymous):

yeah

mathslover (mathslover):

oh k so in the second equation we have : x+y= - 2

OpenStudy (anonymous):

right

mathslover (mathslover):

now put this value of y = x-6 on the place of y in second equation \[\large{x+y=-2}\] \[\large{x+(x-6)=-2}\]

mathslover (mathslover):

got my point now?

OpenStudy (anonymous):

ohh okay i see

mathslover (mathslover):

so we have : \[\large{x+x-6=-2}\] \[\large{2x-6=-2}\] right?

OpenStudy (anonymous):

yeah thats right

mathslover (mathslover):

add 6 both sides what do you get? : \[\large{2x-6+6=-2+6}\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

x=2

mathslover (mathslover):

right

mathslover (mathslover):

good work

OpenStudy (anonymous):

:)) Thanks i actually learned how to do that

mathslover (mathslover):

now we had : y = x-6 put the value of x = 2

mathslover (mathslover):

y = 2-6 = -4

mathslover (mathslover):

No problem and best of luck

OpenStudy (anonymous):

so what will be the final answer?

mathslover (mathslover):

y = -4 and x = 2

OpenStudy (anonymous):

oh yeah thanks again :)

mathslover (mathslover):

np :)